Two mechanics worked on a car. The first mechanic charged 
$ 55 per hour, and the second mechanic charged $110 per hour. The mechanics worked for a combined total of 20 hours, and together they charged a total of $1925. How long did each mechanic work?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39629)

Do they do the work together?
MECHANICS        PAYRATE         TIME        MONEY
Low-priced        55             20-y        55(20-y)
High-priced      110              y         110y
TOTALS                           20          1925

Also notice that one of the mechanics earns at TWICE the rate than the other.

Answer by ikleyn(52873)

.
Two mechanics worked on a car. The first mechanic charged $ 55 per hour,
and the second mechanic charged $110 per hour.
The mechanics worked for a combined total of 20 hours, and together
they charged a total of $1925. How long did each mechanic work?
~~~~~~~~~~~~~~~~

x = hours 2nd mechanic worked, who charged 110 dollars per hour.

20-x = hours 1st mechanic worked, who charged 55 dollars per hour.


The total cost is 110x + 55*(20-x), and the total money equation is THIS

    110x + 55*(20-x) = 1925  dollars.


Simplify it and find x

    110x + 1100 - 55x = 1925

    110 - 55x = 1925 - 1100

       55x    =    825

         x    =    825/55 = 15.


ANSWER.  Second mechanic worked 15 hours.  First mechanic worked (20-15) = 5 hours.


CHECK.  15*110 + 55*5 = 1925,  total money.    ! correct !

Solved.